# Exploring AI-based System Design for Pixel-level Protected Health Information Detection in Medical Images

### v3

@@ 6 Discussion
We investigate four different setups of an end-to-end pipeline for pixel-level PHI detection in medical images. For the localization step, we recommend using a dedicated object detection model instead of relying on the built-in localization component of a generic OCR model (Setup 2) or the localization capability of a multi-modal LLM (Setup 4). Our experimental results demonstrate that separating the localization step and using a localization model trained explicitly on the data reflecting the target domain ensures strong overall performance, even when out-of-the-box OCR tools, like EasyOCR, are used downstream for text extraction (Setup 1). End-to-end GPT-4o (Setup 4) has the drawback of not returning the bounding box coordinates of text regions, which is essential for subsequent PHI redaction. Additionally, imprints are often incompletely identified and hallucinations occur much more frequently compared to other setups, resulting in lower performance. Moreover, there are more pipeline errors related to server issues or unstructured output, which prevents GPT-4o from generating a proper response.
For the text extraction step, we observe that GPT-4o outperforms EasyOCR by having lower word-level and character-level errors. However, when evaluating the end-to-end pipeline performance, the improvement in PHI detection when switching to GPT-4o from EasyOCR (i.e, Setup 3 vs. Setup 1) is minimal. This is because minor OCR errors produced by EasyOCR can easily be circumvented by the robust language model in the text analysis step. While GPT-4o offers superior performance in text extraction, the overall gain in pipeline performance is modest and accompanied by higher latency and increased costs associated with token generation. Therefore, EasyOCR is preferred in our study.
For text analysis, GPT-4o is the default and only choice across four different setups. The model excels at understanding text content and is particularly efficient in identifying PHI and further classifying it into sub-categories. Additionally, it can accurately discern context even when there are slight OCR errors in the extracted text. For instance, Figure 11 illustrates OCR spelling errors related to dates, identifiers, and gender. The model is capable of implicitly correcting these errors and reasoning based on the corrected text. Furthermore, one of the strengths of LLM is its flexibility, which allows for adjustments in instructions through prompts and configurations of output structures. This adaptability makes it easy to redefine PHI and associated tasks across various studies.
@@
A key concern for any AI-based system is its ability to generalize to unseen data. In light of this, we explore the use of pre-trained or foundational models with minimal fine-tuning, applicable solely to the localization model. Evaluation on the hold-out MIDI dataset has demonstrated that the YOLO model, despite being fine-tuned on a different dataset (RadPHI-train), can perform effectively on new datasets containing modalities not encountered during training, including PET, mammography, and ultrasound. Furthermore, the LLM demonstrates adaptability to diverse contexts, as evidenced by our results on the MIDI dataset with modified prompting. This underscores a valuable use case in which users can interactively adjust the prompt to redact or retain specific information in the images, all without the need to develop a new classification model. Overall, these findings emphasize the potential of leveraging existing models to enhance performance across various medical imaging domains.
@@
In terms of limitations, the expansion of GPT-4o’s capabilities beyond basic text analysis to the interpretation of small image crops and, ultimately, entire images reveals interesting findings. While the model demonstrates exceptional performance when analyzing image crops (Setup 3), it encounters a significant increase in technical errors and hallucinations when processing complete images (Table 3). This inefficiency results in incomplete and inaccurate detection of PHI. It also underscores the current limitations of the multimodal language model, particularly in its capacity to handle images with diverse content and to perform complex reasoning tasks that require the integration of information from both images and text. Moreover, the model’s non-deterministic behavior can lead to varying responses for identical inputs. However, by employing structured output formats with Pydantic [27] and setting the temperature parameter to zero, we have observed a marked improvement in response consistency, as indicated by the low standard deviation (Table A1-A4 in Supplementary Material). There remains significant potential to further explore the full capabilities of language models for the PHI detection task. In our future work, we plan to investigate and benchmark various prompting strategies and assess the model’s robustness in languages other than English. Additionally, given the high level of privacy required for medical data, it is essential to either fine-tune a LLM or host it in-house for application development. Smaller LLMs may be particularly interesting to examine due to their ease of deployment and the ability to customize them for specific tasks. Furthermore, we plan to benchmark our pipeline against commercially available products [8, 7, 9, 10] currently available in the market.
@@ Appendix A Results on RadPHI-test
Class Precision Recall F1 Accuracy
Global PHI 0.9988 ± 0.0000 0.9904 ± 0.0016 0.9946 ± 0.0008 0.9908 ± 0.0013
Date 1.0000 ± 0.0000 1.0000 ± 0.0000 1.0000 ± 0.0000 1.0000 ± 0.0000
Patient Name 0.9493 ± 0.0030 0.9961 ± 0.0001 0.9722 ± 0.0016 0.9852 ± 0.0008
Address 0.9976 ± 0.0022 0.9992 ± 0.0018 0.9984 ± 0.0017 0.9992 ± 0.0008
Gender 0.9967 ± 0.0030 0.7414 ± 0.0062 0.8512 ± 0.0059 0.9363 ± 0.0015
Email 1.0000 ± 0.0000 0.9509 ± 0.0033 0.9749 ± 0.0017 0.9864 ± 0.0009
Phone Nr 0.9923 ± 0.0000 0.9923 ± 0.0000 0.9923 ± 0.0000 0.9960 ± 0.0000
Age 1.0000 ± 0.0000 0.9266 ± 0.0051 0.9619 ± 0.0028 0.9812 ± 0.0013
Identifier 0.9948 ± 0.0019 0.9747 ± 0.0030 0.9847 ± 0.0018 0.9928 ± 0.0008
@@
Class Precision Recall F1 Accuracy
Global PHI 0.9875 ± 0.0011 0.9656 ± 0.0013 0.9764 ± 0.0009 0.9604 ± 0.0016
Date 1.0000 ± 0.0000 0.8703 ± 0.0063 0.9307 ± 0.0037 0.9654 ± 0.0017
Patient Name 0.8480 ± 0.0039 0.9308 ± 0.0027 0.8874 ± 0.0023 0.9385 ± 0.0014
Address 0.9812 ± 0.0022 0.9694 ± 0.0036 0.9752 ± 0.0023 0.9878 ± 0.0011
Gender 0.9965 ± 0.0032 0.6934 ± 0.0034 0.8178 ± 0.0029 0.9245 ± 0.0012
Email 0.9991 ± 0.0019 0.8356 ± 0.0042 0.9101 ± 0.0031 0.9542 ± 0.0015
Phone Nr 0.9909 ± 0.0032 0.8339 ± 0.0032 0.9056 ± 0.0027 0.9548 ± 0.0013
Age 0.9929 ± 0.0026 0.6547 ± 0.0045 0.7891 ± 0.0032 0.9103 ± 0.0010
Identifier 0.8492 ± 0.0061 0.8886 ± 0.0056 0.8685 ± 0.0038 0.9361 ± 0.0019
@@
Class Precision Recall F1 Accuracy
Global PHI 0.9995 ± 0.0007 0.9972 ± 0.0006 0.9983 ± 0.0005 0.9972 ± 0.0008
Date 1.0000 ± 0.0000 1.0000 ± 0.0000 1.0000 ± 0.0000 1.0000 ± 0.0000
Patient Name 0.9908 ± 0.0021 0.9969 ± 0.0017 0.9938 ± 0.0008 0.9968 ± 0.0004
Address 1.0000 ± 0.0000 0.9959 ± 0.0001 0.9980 ± 0.0000 0.9990 ± 0.0000
Gender 0.9957 ± 0.0000 0.9483 ± 0.0063 0.9714 ± 0.0033 0.9864 ± 0.0016
Email 1.0000 ± 0.0000 0.9964 ± 0.0000 0.9985 ± 0.0007 0.9990 ± 0.0000
Phone Nr 0.9985 ± 0.0021 0.9977 ± 0.0021 0.9981 ± 0.0014 0.9990 ± 0.0007
Age 1.0000 ± 0.0000 0.9913 ± 0.0018 0.9957 ± 0.0009 0.9978 ± 0.0004
Identifier 1.0000 ± 0.0000 0.9797 ± 0.0082 0.9897 ± 0.0042 0.9952 ± 0.0019
@@
Class Precision Recall F1 Accuracy
Global PHI 0.9929 ± 0.0015 0.9482 ± 0.0030 0.9701 ± 0.0013 0.9506 ± 0.0021
Date 0.9410 ± 0.0072 0.9513 ± 0.0050 0.9461 ± 0.0030 0.9722 ± 0.0017
Patient Name 0.6913 ± 0.0147 0.9025 ± 0.0146 0.7827 ± 0.0082 0.8786 ± 0.0063
Address 0.8725 ± 0.0048 0.9571 ± 0.0083 0.9128 ± 0.0039 0.9563 ± 0.0021
Gender 0.9160 ± 0.0090 0.7560 ± 0.0095 0.8284 ± 0.0081 0.9278 ± 0.0024
Email 0.9813 ± 0.0019 0.9034 ± 0.0075 0.9407 ± 0.0038 0.9696 ± 0.0017
Phone Nr 0.9648 ± 0.0063 0.8697 ± 0.0117 0.9148 ± 0.0068 0.9597 ± 0.0031
Age 0.9246 ± 0.0089 0.9431 ± 0.0067 0.9337 ± 0.0038 0.9666 ± 0.0022
Identifier 0.8026 ± 0.0109 0.9550 ± 0.0073 0.8722 ± 0.0084 0.9359 ± 0.0050
@@ Appendix B Results on MIDI
Class Precision Recall F1 Accuracy
Global PHI 0.9900 ± 0.0034 0.9962 ± 0.0035 0.9931 ± 0.0014 0.9890 ± 0.0022
Date 1.0000 ± 0.0000 0.9948 ± 0.0077 0.9974 ± 0.0039 0.9970 ± 0.0045
Patient Name 0.9735 ± 0.0076 0.9885 ± 0.0105 0.9809 ± 0.0059 0.9800 ± 0.0061
Address 0.9631 ± 0.0291 0.9889 ± 0.0152 0.9755 ± 0.0110 0.9910 ± 0.0042
Gender 0.9965 ± 0.0079 0.9492 ± 0.0000 0.9722 ± 0.0038 0.9840 ± 0.0022
Phone Nr 0.9723 ± 0.0253 0.9619 ± 0.0399 0.9663 ± 0.0138 0.9930 ± 0.0027
Age 0.9637 ± 0.0133 0.9606 ± 0.0083 0.9621 ± 0.0075 0.9750 ± 0.0050
Identifier 0.7622 ± 0.0093 0.9899 ± 0.0106 0.8612 ± 0.0088 0.8740 ± 0.0082
@@
Class Precision Recall F1 Accuracy
Global PHI 0.9987 ± 0.0028 0.9924 ± 0.0029 0.9955 ± 0.0017 0.9930 ± 0.0027
Date 0.9983 ± 0.0038 0.9931 ± 0.0038 0.9957 ± 0.0030 0.9950 ± 0.0035
Patient Name 0.9923 ± 0.0124 0.9808 ± 0.0000 0.9865 ± 0.0062 0.9860 ± 0.0065
Address 0.9472 ± 0.0255 0.9889 ± 0.0152 0.9675 ± 0.0181 0.9880 ± 0.0067
Gender 0.9965 ± 0.0078 0.9627 ± 0.0142 0.9793 ± 0.0099 0.9880 ± 0.0057
Phone Nr 0.9735 ± 0.0391 1.0000 ± 0.0000 0.9862 ± 0.0204 0.9970 ± 0.0045
Age 0.9639 ± 0.0079 0.9697 ± 0.0000 0.9668 ± 0.0040 0.9780 ± 0.0027
Identifier 0.9920 ± 0.0118 0.9342 ± 0.0165 0.9621 ± 0.0087 0.9709 ± 0.0066
@@ Appendix D Examples of pipeline output
Figure D1: Comparison of outputs from four setups: Setups 1, 3, and 4 demonstrate similar response patterns, with Setup 3 accurately identifying the gender. In contrast, Setup 2 exhibits incomplete text extraction, missing the patient’s name and gender.
Figure D2: Comparison of outputs from four setups: Setup 1 and Setup 3 accurately classify PHI categories despite overlapping address image regions. Setup 2 fails to recognize the age class, while Setup 4 illustrates model hallucination, generating responses that are not based on the provided input.

### v4

@@ 6 Discussion
Text localization
For the localization step, we recommend using a dedicated object detection model instead of relying on the built-in localization component of a generic OCR model (Setup 2) or the localization capability of a multi-modal LLM (Setup 4). Our experimental results demonstrate that separating the localization step and using a localization model that can generalize well across different medical imaging modalities ensures strong overall performance, even when out-of-the-box OCR tools, like EasyOCR, are used downstream for text extraction (Setup 1). End-to-end GPT-4o (Setup 4) has the drawback of not returning the bounding box coordinates of text regions, which are essential for subsequent PHI redaction. Additionally, imprints are often incompletely identified and hallucinations occur much more frequently compared to other setups, resulting in lower performance. Moreover, in this setup, more pipeline errors related to server issues or unstructured output occur, which prevents GPT-4o from generating proper responses.
Text extraction
When evaluating the OCR performance separately, GPT-4o outperforms EasyOCR by having lower word-level and character-level errors. However, for the end-to-end pipeline performance, the improvement in PHI detection when switching to GPT-4o from EasyOCR (i.e, Setup 3 vs. Setup 1) is minimal. This is because minor OCR errors produced by EasyOCR can easily be compensated by the robust language model in the text analysis step. While GPT-4o offers superior performance in text extraction, the overall gain in pipeline performance is modest and accompanied by higher latency and increased costs associated with token generation. Therefore, EasyOCR is preferred in our study.
Text analysis
For text analysis, GPT-4o is the default and only choice across all four different setups. The model excels at understanding text content and is particularly efficient in identifying PHI and further classifying it into sub-categories. Additionally, it can accurately discern context even when there are slight OCR errors in the extracted text. For instance, Figure 10 illustrates OCR errors related to dates, identifiers, and names. The model is capable of implicitly correcting these errors and reasoning based on the corrected text. Furthermore, an outstanding strength of LLM-based text analysis is the possibility to easily configure rules, sequential steps, and the output structure of the analysis through prompts. This adaptability makes it easy to redefine PHI and associated tasks across various studies or to accommodate different guidelines.
@@
Generalization
A key concern for any AI-based system is its ability to generalize to unseen data. In context of PHI detection, we believe the main source of variability stems from the text representation such as font size, color, and background, as well as the content itself, which may include PHI in various languages or formats not explicitly defined within the default PHI criteria. We address this by leveraging pretrained models whenever possible, requiring minimal fine-tuning. By choosing pretrained models, we preserve the models’ learned abstractions and only guide their reasoning towards specific criteria. For example, in our optimal setup (Setup 1), the only model that requires fine-tuning is the YOLOv11 model for imprint localization. By covering key imaging modalities, generating diverse imprints and utilizing numerous data augmentations methods during training, we achieve during inference that the YOLOv11 model performs robustly in unseen modalities that are not encountered during training, including PET, mammography, and ultrasound as part of the MIDI dataset. The other two models, namely EasyOCR and GPT-4o, work robustly with either no fine-tuning or only minor adjustments in the case of GPT-4o. On the MIDI dataset, which we have little control over the format of the imprints that appear in the images exported from the viewer. The results show that even without optimizing the prompt to tailor the formatting or representation, we achieve a decent recall rate of 0.9639, missing on average 19.8 out of 549 PHI instances. The precision is low, 0.8797, with 72.4 false positives on average due to unadjusted guidelines for classification. However, after adjusting the prompt to define specific criteria for inclusion and exclusion, we reduce the false positives to 15.4, improving the precision to 0.9721. This adaptation occurs only in the text analysis to conform to the new guidelines we select, while the text localization and extraction processes continue to perform well, even with the new dataset. While it is true that the LLM may struggle with new guidelines not covered in the initial prompt, modifying the PHI classification criteria is relatively straightforward using natural language. This stands in contrast with the complex engineering needed to customize rule-based classification systems or to fine-tune models with fixed rules. Additionally, it is less complicated to extend PHI classification to encompass diverse languages with LLMs, which is an advantage of using LLMs for this purpose.
Capabilities of multi-modal LLM
@@
The expansion of GPT-4o’s capabilities beyond basic text analysis to the interpretation of small image crops and, ultimately, entire images reveals interesting findings. While the model demonstrates exceptional performance when analyzing image crops (Setup 3), it encounters a significant increase in technical errors and hallucinations when processing complete images in Setup 4 (Table 6). This inefficiency results in incomplete and inaccurate detection of PHI. It also underscores the current limitations of the multimodal language model, particularly in its capacity to handle images with diverse content and to perform complex reasoning tasks that require the integration of information from both images and text. Moreover, the model’s non-deterministic behavior can lead to varying responses for identical inputs. However, by employing structured output formats with Pydantic [33] and setting the temperature parameter to zero, we have observed a marked improvement in response consistency, as indicated by the low standard deviation (Table A1-A4 in Supplementary Material). There remains significant potential to further explore the full capabilities of language models for the PHI detection task. In our future work, we plan to investigate and benchmark various prompting strategies and assess the model’s robustness in languages other than English. Additionally, given the high level of privacy required for medical data, it is essential to either fine-tune a LLM or host it in-house. Smaller LLM architectures are here of particular interest due to their ease of deployment, increased inference speed, reduced costs and the ability to customize them for specific tasks.
Limitations and future work
In our study, we focus on identifying the most effective setup for constructing a PHI pipeline, without delving deeply into benchmarking specific model types or variants to determine which offers the best performance. By optimizing the models used in each module separately, we acknowledge that our findings may not necessarily apply to particular models. For instance, we believe that some commercial OCR tools can process images effectively without requiring further fine-tuning, suggesting that Setup 2 may perform comparably to Setup 1. Moreover, our evaluation does not include testing our proposed approaches on real-world data containing PHI. In general, it is challenging to obtain medical image data containing actual PHI imprints for research and analysis in a privacy compliant way. Regulatory guidelines require that detected PHI is removed or de-identified by the data processor. In view of this, the MIDI dataset is created by utilizing an industry-standard medical image viewer to overlay synthetically generated PHI on medical images. Exporting image data directly from this viewer allows us to generate test data that is very close to reality. We recognize that real-world datasets are very heterogeneous and may comprise a diverse range of modalities and scanners. Hence, we plan to further extend our benchmark in future work to incorporate a wider variety of model types and curate additional datasets for validation, and extend the language beyond English.
@@ Appendix A Results on RadPHI-test
Class Case Precision Case Recall Instance Precision Instance Recall
PHI Presence 1.0000 ± 0.0000 0.9923 ± 0.0000 0.9869 ± 0.0008 0.9808 ± 0.0012
Date 1.0000 ± 0.0000 0.9963 ± 0.0000 0.9963 ± 0.0000 0.9963 ± 0.0000
Patient Name 0.9778 ± 0.0032 0.9815 ± 0.0017 0.9642 ± 0.0033 0.9731 ± 0.0000
Address 0.9952 ± 0.0018 1.0000 ± 0.0000 0.9857 ± 0.0045 1.0000 ± 0.0000
Email 1.0000 ± 0.0000 0.9496 ± 0.0000 1.0000 ± 0.0000 0.9496 ± 0.0000
Phone Nr 1.0000 ± 0.0000 0.9908 ± 0.0021 1.0000 ± 0.0000 0.9908 ± 0.0021
Identifier 0.9916 ± 0.0000 0.9949 ± 0.0035 0.9815 ± 0.0023 0.9848 ± 0.0023
@@
Class Case Precision Case Recall
PHI Presence 0.9852 ± 0.0007 0.9743 ± 0.0016
Date 0.9983 ± 0.0023 0.8689 ± 0.0065
Patient Name 0.8547 ± 0.0011 0.9323 ± 0.0058
Address 0.9942 ± 0.0022 0.9750 ± 0.0018
Email 1.0000 ± 0.0000 0.8489 ± 0.0025
Phone Nr 0.9908 ± 0.0000 0.8254 ± 0.0021
Identifier 0.8301 ± 0.0027 0.9114 ± 0.0000
@@
Class Case Precision Case Recall Instance Precision Instance Recall
PHI Presence 0.9954 ± 0.0007 0.9995 ± 0.0007 0.9859 ± 0.0006 0.9778 ± 0.0021
Date 1.0000 ± 0.0000 0.9655 ± 0.0041 1.0000 ± 0.0000 0.9655 ± 0.0041
Patient Name 1.0000 ± 0.0000 0.9923 ± 0.0047 0.9915 ± 0.0017 0.9838 ± 0.0032
Address 0.9959 ± 0.0029 0.9798 ± 0.0040 0.9959 ± 0.0029 0.9798 ± 0.0040
Email 1.0000 ± 0.0000 0.9849 ± 0.0053 1.0000 ± 0.0000 0.9849 ± 0.0053
Phone Nr 0.9961 ± 0.0000 0.9823 ± 0.0034 0.9961 ± 0.0000 0.9823 ± 0.0034
Identifier 1.0000 ± 0.0000 0.9755 ± 0.0046 0.9913 ± 0.0000 0.9671 ± 0.0046
@@
Class Case Precision Case Recall
PHI Presence 0.9918 ± 0.0023 0.9017 ± 0.0066
Date 0.9309 ± 0.0087 0.8966 ± 0.0146
Patient Name 0.7532 ± 0.0084 0.8469 ± 0.0202
Address 0.9256 ± 0.0184 0.8855 ± 0.0175
Email 0.9950 ± 0.0054 0.8633 ± 0.0171
Phone Nr 0.9843 ± 0.0090 0.7631 ± 0.0169
Identifier 0.8809 ± 0.0230 0.8675 ± 0.0238
@@ Appendix B Results on MIDI
Class Case Precision Case Recall Instance Precision Instance Recall
PHI Presence 0.9509 ± 0.0127 0.9986 ± 0.0030 0.8797 ± 0.0096 0.9639 ± 0.0047
Date 0.9931 ± 0.0039 0.9914 ± 0.0000 0.9553 ± 0.0081 0.9082 ± 0.0048
Patient Name 0.9962 ± 0.0052 0.9981 ± 0.0043 0.9851 ± 0.0055 0.9863 ± 0.0081
Address 0.8931 ± 0.0305 0.9222 ± 0.0232 0.9851 ± 0.0055 0.9863 ± 0.0081
Phone Nr 0.9193 ± 0.0186 0.9714 ± 0.0261 0.9073 ± 0.0189 0.9750 ± 0.0228
Identifier 0.7427 ± 0.0188 0.9772 ± 0.0057 0.6589 ± 0.0167 0.9746 ± 0.0060
@@
Class Case Precision Case Recall Instance Precision Instance Recall
PHI Presence 0.9697 ± 0.0035 1.0000 ± 0.0000 0.9721 ± 0.0032 0.9781 ± 0.0043
Date 1.0000 ± 0.0000 0.9914 ± 0.0000 0.9893 ± 0.0064 0.9758 ± 0.0076
Patient Name 1.0000 ± 0.0000 0.9769 ± 0.0129 0.9776 ± 0.0035 0.9752 ± 0.0116
Address 0.9500 ± 0.0116 0.9444 ± 0.0278 0.9070 ± 0.0092 0.9487 ± 0.0256
Phone Nr 1.0000 ± 0.0000 0.9143 ± 0.0213 0.9648 ± 0.0197 0.9083 ± 0.0186
Identifier 0.9398 ± 0.0136 0.9848 ± 0.0057 0.9325 ± 0.0115 0.9814 ± 0.0038
@@ Appendix D Examples of pipeline output
Figure D1: Comparison of outputs from four setups: Setups 1, 3, and 4 demonstrate similar response patterns. In contrast, Setup 2 exhibits duplicate text extraction, resulting in similar entries.
Figure D2: Comparison of outputs from four setups: Setup 1, 2, and 3 accurately classify PHI category while Setup 4 illustrates model hallucination, generating responses that are not based on the provided input.